OpenStudy (anonymous):

Find the distance for the following points A ( -2, -3) B (-7, -8). If necessary round your answer to the nearest tenth.

OpenStudy (anonymous):

\[\sqrt{25+25}=\sqrt{50}\]

OpenStudy (anonymous):

11 2) 7.1 3) 7 4) -7.1